# Leveraging NTPs for Efficient Hallucination Detection in VLMs

### v1

@@ Abstract
Ofir Azachithanks: Equal contribution.,1, Kfir Eliyahu11footnotemark: 1,1, Eyal El Ani11footnotemark: 1,1, Rom Himelstein11footnotemark: 1,1, Roi Reichart1, Yuval Pinter2, Nitay Calderon1 1Department of Data and Decision Science, Technion - Israel Institute of Technology, 2Department of Computer Science, Ben-Gurion University of the Negev Correspondence: romh@campus.technion.ac.il.
@@ 1 Introduction
A growing body of research shows that VLMs often rely heavily on linguistic priors (Zhu et al., 2024; Guan et al., 2024; eun Cho and Maeng, 2025; Wang et al., 2024), and may even prioritize them over conflicting visual evidence (Luo et al., 2024; Wu et al., 2024). These findings suggest that hallucinations in VLMs may stem, at least in part, from biases in their language modeling components, rather than solely from limitations in visual understanding. Based on these insights, we introduce a novel dataset specifically curated to examine the relationship between NTPs and hallucinations in VLMs. We believe this dataset will serve as a valuable resource for future research in this area. [^1] Code and dataset will be released upon publication. Using this dataset, we evaluate the effectiveness of NTPs generated by LLaVA-1.5 and LLaVA-1.6 (Liu et al., 2024) for hallucination detection. As baselines, we include predictions from both LLaVA and PaliGemma (Beyer et al., 2024), and also use these predictions as additional input features to traditional ML models.
@@ 5 Experimental Setup
VLM Predictors To evaluate the effectiveness of probe-based hallucination detection, we employ two VLM predictors. The first is a LLaVA-based predictor,555huggingface.co/llava-hf/llava-1.5-7b-hf; huggingface.co/llava-hf/llava-v1.6-mistral-7b-hf corresponding to the same VLM that generated the image description. The rationale is to compare the performance of traditional ML models that rely on the VLM’s NTPs with that of using the same VLM for self-verification of its own generated content. The second VLM predictor is an external model, PaliGemma.666huggingface.co/google/PaliGemma-3b-pt-224 Naturally, using an external VLM also imposes additional computational and memory overhead.

### v2

@@ Abstract
Ofir Azachithanks: Equal contribution.,1, Kfir Eliyahu11footnotemark: 1,1, Eyal El Ani11footnotemark: 1,1, Rom Himelstein11footnotemark: 1,1, Roi Reichart1, Yuval Pinter2, Nitay Calderon1 1Department of Data and Decision Science, Technion - Israel Institute of Technology, 2 Faculty of Computer and Information Science, Ben-Gurion University of the Negev Correspondence: romh@campus.technion.ac.il.
@@ 1 Introduction
A growing body of research shows that VLMs often rely heavily on linguistic priors (Zhu et al., 2024; Guan et al., 2024; eun Cho and Maeng, 2025; Wang et al., 2024), and may even prioritize them over conflicting visual evidence (Luo et al., 2024; Wu et al., 2024). These findings suggest that hallucinations in VLMs may stem, at least in part, from biases in their language modeling components, rather than solely from limitations in visual understanding. Based on these insights, we introduce a novel dataset specifically curated to examine the relationship between NTPs and hallucinations in VLMs. We believe this dataset will serve as a valuable resource for future research in this area. Using this dataset, we evaluate the effectiveness of NTPs generated by LLaVA-1.5 and LLaVA-1.6 (Liu et al., 2024) for hallucination detection. As baselines, we include predictions from both LLaVA and PaliGemma (Beyer et al., 2024), and also use these predictions as additional input features to traditional ML models.
@@ 5 Experimental Setup
VLM Predictors To evaluate the effectiveness of probe-based hallucination detection, we employ two VLM predictors. The first is a LLaVA-based predictor,444huggingface.co/llava-hf/llava-1.5-7b-hf; huggingface.co/llava-hf/llava-v1.6-mistral-7b-hf corresponding to the same VLM that generated the image description. The rationale is to compare the performance of traditional ML models that rely on the VLM’s NTPs with that of using the same VLM for self-verification of its own generated content. The second VLM predictor is an external model, PaliGemma.555huggingface.co/google/PaliGemma-3b-pt-224 Naturally, using an external VLM also imposes additional computational and memory overhead.